A sailboat travels 20 miles downstream in 3 hours. It returns in 4 hours. Find the speed of the sailboat in still water and the rate of the current.
Help, no idea.
Thanks.

Answer by stanbon(48558)

A sailboat travels 20 miles downstream in 3 hours. It returns in 4 hours. Find the speed of the sailboat in still water and the rate of the current.
----------------------
Downstream DATA:
distance = 20 miles; time = 3 hrs. ;rate = d/t = 20/3 mph
---------------------------------------
Upstream DATA:
distance = 20 miles; time = 4 hrs. ;rate = d/t = 20/4 = 5 mph
-----------------------------------------
Let still water speed be "b"; Let current speed be "c"
Downstream: b + c = 20/3 mph
Upstream : b - c = 5
----------------------
2b = 35/3
b = 35/6 = 5 4/6 mph (boat speed in still water)
-------
Use b -c = 5 to solve for "c".
35/6 - c = 30/6
c = 5/6 mph (current speed)
================================
Cheers,
Stan H.